# Equivariant Neural Networks for General Linear Symmetries on Lie Algebras

### v1

@@ B.1 Proof of Non-degeneracy and Ad-invariance of Modified Killing Form $B_{e}$
Definition B.1 (Reductive decomposition).
@@
Definition B.2 (Modified Killing form on a reductive Lie algebra).
@@
Proposition B.1 (Block–orthogonality and restrictions).
@@
Proposition B.3 ($\mathrm{Ad}$–invariance on the identity component).
@@
The restriction to $\mathfrak{g},\mathfrak{g}]$ equals $B$, which is $\mathrm{ad}$–invariant. If $Z\in\mathfrak{z}(\mathfrak{g})$ then $[X,Z]=0$ for all $X$, so any bilinear form on $\mathfrak{z}(\mathfrak{g})$ is automatically $\mathrm{ad}$–invariant. Using Proposition B.1 and bilinearity gives the displayed identity. Equivalence with $\mathrm{Ad}$–invariance on $G^{\circ}$ follows by integrating the infinitesimal relation along paths in $G^{\circ}$. ∎
Remark 2 (Invariance for nonconnected groups).
@@
Replacing $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ by $\langle\cdot,\cdot\rangle_{\mathfrak{z}}^{\mathrm{avg}}$ in 18 makes $\widetilde{B}$ invariant under all of $G$.
Appendix C Connections to existing bilinear forms
We demonstrate how the trace-based form $\widetilde{B}$ (Eq. 2) unifies and recovers prior bilinear constructions in the regimes used by Lie Neurons (Lin et al., 2024a) and Vector Neurons (Deng et al., 2021) . The discussion below states precise conditions under which $\widetilde{B}$ (i) equals the Killing-form contractions on semisimple inputs and (ii) is proportional to the Vector Neuron inner product under the $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$ isomorphism. Our single form applies on the full reductive algebra $\mathfrak{gl}(n)$ including these specialized approaches, which encompasses general $n\times n$ matrix-valued inputs.
C.1 Reduction to the Killing form on the semisimple ideal
Write $X=X_{0}+\tfrac{1}{n}\mathrm{tr}(X)I$ with $X_{0}\in\mathfrak{sl}(n)$. Using Eq. 2 we obtain
$\widetilde{B}(X,Y)=2n\,\mathrm{tr}(X_{0}Y_{0})+n^{2}\!\left(\tfrac{1}{n}\mathrm{tr}X\right)\!\left(\tfrac{1}{n}\mathrm{tr}Y\right).$ (22)
Hence, when inputs are restricted to the semisimple ideal $\mathfrak{g},\mathfrak{g}]=\mathfrak{sl}(n)$ (so $\mathrm{tr}(X)=\mathrm{tr}(Y)=0$), the center contribution vanishes and
$\widetilde{B}|_{\mathfrak{sl}(n)}(X,Y)=2n\cdot\mathrm{tr}(XY)=B_{\mathfrak{sl}(n)}(X,Y),$ (23)
i.e. $\widetilde{B}$ coincides with the (scaled) Killing form used in Lie Neurons. More generally, for any semisimple subalgebra $\mathfrak{h}\subset[\mathfrak{g},\mathfrak{g}]$, the restriction $\widetilde{B}|_{\mathfrak{h}}$ matches the Killing-form contraction on $\mathfrak{h}$ up to global scaling.
C.2 $\mathfrak{so}(3)$ example: recovery of the Vector Neuron inner product
Recall the hat map $\widehat{(\cdot)}:\mathbb{R}^{3}\to\mathfrak{so}(3)$ with
$\widehat{v}=\begin{bmatrix}0&-v_{3}&v_{2}\\ v_{3}&0&-v_{1}\\ -v_{2}&v_{1}&0\end{bmatrix},\qquad\mathrm{tr}(\widehat{v}\,\widehat{w})=-2\,v^{\top}w.$ (24)
For $X=\widehat{v}$ and $Y=\widehat{w}$, Eq. 2 gives
$\widetilde{B}(X,Y)=2n\cdot\mathrm{tr}(XY)=2n\cdot(-2\,v^{\top}w)=(-4n)\,v^{\top}w,$ (25)
i.e. $\widetilde{B}$ is proportional to the Euclidean inner product on $\mathbb{R}^{3}$. The proportionality constant depends only on $n$; in practice this constant is absorbed by adjacent learnable linear layers or normalization, yielding behaviour identical to the inner product used in Vector Neurons to harmless scaling.
These remarks justify using the single, nondegenerate $\widetilde{B}$ across heterogeneous input types while preserving compatibility with prior architectures.
Appendix D Detailed Layer Formulations
@@
This section provides the precise mathematical definitions and equivariance proofs for the core components of the ReLN architecture. We consider the input to a layer as a tensor $x\in\mathbb{R}^{K\times C}$, representing $C$ feature channels where $K=\dim\mathfrak{g}$. Each column $x_{c}\in\mathbb{R}^{K}$ is the vector representation of a feature. We use the wedge ($\wedge$) and vee ($\vee$) operators to map between the vector form $x_{c}$ and the matrix form $X_{c}\in\mathfrak{g}$.
D.1 Equivariant Linear Layer
@@
The ReLN-Linear layer applies a linear map to the channel dimension of the input tensor $x$:
$f_{\mathrm{ReLN-Lin}}(x;W)=xW,\quad\text{where }W\in\mathbb{R}^{C\times C^{\prime}}.$ (26)
@@ D.2 Equivariant Nonlinearities
This layer rectifies a feature based on its alignment with a learnable direction. Given the input tensor $x$, we first compute per-channel vector directions $d=xU$. The nonlinearity for the input $x$ is then defined as:
$f_{\mathrm{ReLN-ReLU}}(x)=\begin{cases}x,&\text{if }\widetilde{B}(x^{\wedge},d^{\wedge})\leq 0,\\ x+\widetilde{B}(x^{\wedge},d^{\wedge})d,&\text{otherwise}.\end{cases}$ (28)
Since all operations—the linear map to compute $d_{i}$, the bilinear form $\widetilde{B}$, and vector addition/scaling—are equivariant, the entire function is equivariant. The leaky variant $f_{\mathrm{ReLN-LeakyReLU}}(x)=\alpha\,x+(1-\alpha)\,f_{\mathrm{ReLN-ReLU}}(x)$ follows directly.
@@
This layer uses the matrix commutator, a natural $\mathrm{Ad}$-equivariant primitive, to create learnable interactions between channels. Let the input be a batch of features represented by their vector coordinates, $x\in\mathbb{R}^{B\times C_{\mathrm{in}}\times K}$. First, two independent linear maps (with learnable weights $W_{a},W_{b}\in\mathbb{R}^{C_{\mathrm{in}}\times C_{\mathrm{out}}}$) transform the input features into two intermediate tensors, $u,v\in\mathbb{R}^{B\times C_{\mathrm{out}}\times K}$. The Lie bracket is then computed channel-wise between the corresponding feature vectors of $u$ and $v$. This produces an update tensor, $\Delta x\in\mathbb{R}^{B\times C_{\mathrm{out}}\times K}$, where each vector is defined as:
$(\Delta x)_{b,c^{\prime},:}=\left[(u_{b,c^{\prime},:})^{\wedge},(v_{b,c^{\prime},:})^{\wedge}\right]^{\vee}.$ (29)
This update is added to the input for a residual connection (requiring $C_{\mathrm{in}}=C_{\mathrm{out}}$ for the shapes to match):
@@
Each step in this process (linear map, the Lie bracket, and vee/hat operations) is equivariant under the adjoint action, making the entire block equivariant $f_{\mathrm{ReLN-Bracket}}(\mathrm{Ad}_{g}x)=\mathrm{Ad}_{g}(f_{\mathrm{ReLN-Bracket}}(x))$ for all $g\in G$.
Appendix E Experimental Details
@@
E.1 Model Architectures and Implementation Details
@@
For the Top-Tagging task, our model is a modification of the LorentzNet architecture (Gong et al., 2022). We adapt its Lorentz Group Equivariant Blocks (LGEBs) by replacing the invariant feature computation with our proposed bilinear form. A detailed description of the architecture, our modifications, and training protocol is provided in Appendix F.
@@
In this task, we compare our ReLN model against two baseline families: a non-equivariant 1D ResNet and an equivariant Vector Neurons (VN) model. The specific implementation details and architectural choices for each model are provided next in Appendix G.
E.2 Platonic Solid Classification on $\mathfrak{sl}(3)$
@@
Appendix F Top Tagging Experiment: Framework, Proof, and Implementation
This appendix provides the complete details for our jet tagging experiment. We first present the geometric framework and the mathematical proof of our Lorentz-equivariant embedding, and then describe the model architecture and training protocol.
F.1 Group Action Equivariance via Embedding Map
To process four-momenta within our Lie-algebraic framework, we require an embedding that translates the action of the Lorentz group into an adjoint action on a matrix space. This is achieved by lifting the four-vector into $\mathfrak{gl}(5)$.
Definition F.1 (Lorentz-Compatible Embedding).
Given a four-vector $p\in\mathbb{R}^{4}$ and the Minkowski metric $\eta=\operatorname{diag}(1,-1,-1,-1)$, we define its embedding $\varphi(p)$ into $\mathfrak{gl}(5)$ as:

### v2

@@ C.1 Proof of Non-degeneracy and Ad-invariance of Modified Killing Form $B_{e}$
Definition C.1 (Reductive decomposition).
@@
Definition C.2 (Modified Killing form on a reductive Lie algebra).
@@
Proposition C.4 (Block–orthogonality and restrictions).
@@
Proposition C.6 ($\mathrm{Ad}$–invariance on the identity component).
@@
The restriction to $\mathfrak{g},\mathfrak{g}]$ equals $B$, which is $\mathrm{ad}$–invariant. If $Z\in\mathfrak{z}(\mathfrak{g})$ then $[X,Z]=0$ for all $X$, so any bilinear form on $\mathfrak{z}(\mathfrak{g})$ is automatically $\mathrm{ad}$–invariant. Using Proposition C.4 and bilinearity gives the displayed identity. Equivalence with $\mathrm{Ad}$–invariance on $G^{\circ}$ follows by integrating the infinitesimal relation along paths in $G^{\circ}$. ∎
Remark C.7 (Invariance for nonconnected groups).
@@
Replacing $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ by $\langle\cdot,\cdot\rangle_{\mathfrak{z}}^{\mathrm{avg}}$ in Equation 31 makes $\widetilde{B}$ invariant under all of $G$.
Appendix D Summary of Key Lie Groups and Algebras
To provide a comprehensive overview of the geometric structures discussed in this work, Table 7 summarizes the Lie groups $G$, their associated Lie algebras $\mathfrak{g}$, and their application domains. We categorize each group by its algebraic properties—specifically distinguishing between reductive, semisimple, and compact types—and detail their specific roles within both the broader literature and our experimental framework. This unified taxonomy situates ReLNs as a general-purpose backbone capable of handling a wide array of linear symmetries encountered in scientific machine learning.
Table 7: Comprehensive survey of Lie groups, algebras, and their applications. We highlight the algebraic classifications that dictate the choice of invariant bilinear forms and contrast general literature examples with our specific experimental tasks.
Group ($G$) Algebra ($\mathfrak{g}$) Algebraic Type General Applications & Refs Relevance to ReLN Framework
General Linear $\mathfrak{gl}(n)$ Reductive General linear transformations, stress-strain tensors (Basu et al., 2025; Finzi et al., 2021). Core Domain. Primary target of ReLNs; resolves the Killing form degeneracy on $\mathfrak{gl}(n)$.
Special Linear $\mathfrak{sl}(3)$ Semisimple Homography classification, 3D vision, SLAM (Lin et al., 2024a; Finzi et al., 2021). Algebraic Benchmark. Validates generalization on semisimple non-compact subgroups (Section  5.1.1).
Symplectic $\mathfrak{sp}(4)$ Semisimple Hamiltonian mechanics, phase space dynamics (Lin et al., 2024a; Finzi et al., 2021). Algebraic Benchmark. Tests the ability to learn highly nonlinear invariants on conserved systems (Section  5.1.2).
Special Orthogonal $\mathfrak{so}(3)$ Compact 3D Point clouds, state estimation, rigid body rotations (Deng et al., 2021; Son et al., 2024; Thomas et al., 2018). Robotics Application. Lifts 3D velocity vectors into $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$ (Sections 5.2 and 5.3).
Lorentz Group $\mathfrak{so}(1,3)^{+}$ Semisimple Particle physics, jet tagging, relativistic collisions (Bogatskiy et al., 2020; Batatia et al., 2023). Physics Application. 4-momenta embedded into $\mathfrak{gl}(5)$ to unify left-action with adjoint action (Section  H.1.
Special Unitary $\mathfrak{su}(3)$ Compact Quantum Chromodynamics (QCD), lattice gauge theory (Favoni et al., 2022). General Example. Illustrates framework applicability to complex-valued compact groups.
SPD Manifold $\mathrm{SPD}(n)$ Riemannian Geometric uncertainty, inertia tensors, diffusion MRI (Ma et al., 2025; Magnus, 1985). Geometric Uncertainty. Covariances mapped to $\mathfrak{gl}(n)$ (Sections 5.2 and 5.3).
Note: $\mathrm{SPD}(n)$ is not a group but a manifold representable as the quotient space $\mathrm{GL}(n)/\mathrm{O}(n)$. We process it by mapping to the linear subspace $\mathfrak{sym}(n)\subset\mathfrak{gl}(n)$ via the matrix logarithm (Ma et al., 2025).
Appendix E Detailed Layer Formulations
@@
This section provides the precise mathematical definitions and equivariance proofs for the ReLN architecture. To maintain consistency with the main text, we denote the input tensor as $x\in\mathbb{R}^{B\times K\times C}$, where $B$ is the batch size, $K=\dim\mathfrak{g}$ is the geometric dimension, and $C$ is the number of feature channels. Each column $x_{b,:,c}\in\mathbb{R}^{K}$ corresponds to a matrix $X_{b,c}\in\mathfrak{g}$ via the vee/hat isomorphism.
E.1 Equivariant Linear Layer
@@
The ReLN-Linear layer applies a linear map to the channel dimension (the last axis). For an input $x\in\mathbb{R}^{B\times K\times C}$ and weights $W\in\mathbb{R}^{C\times C^{\prime}}$, the operation is defined as:
$f_{\mathrm{ReLN-Lin}}(x;W)=xW\in\mathbb{R}^{B\times K\times C^{\prime}}.$ (35)
@@ E.2 Equivariant Nonlinearities
This layer rectifies a feature based on its alignment with a learnable direction. Given the input $x\in\mathbb{R}^{B\times K\times C}$, we first compute equivariant directions $d=f_{\mathrm{ReLN-Lin}}(x;U)$ where $U\in\mathbb{R}^{C\times C}$. The nonlinearity is then applied channel-wise as a scalar-gated update:
$f_{\mathrm{ReLN-ReLU}}(x)_{b,:,c}=\begin{cases}x_{b,:,c},&\text{if }\widetilde{B}(x_{b,:,c}^{\wedge},d_{b,:,c}^{\wedge})\leq 0,\\ x_{b,:,c}+\sigma\big(\widetilde{B}(x_{b,:,c}^{\wedge},d_{b,:,c}^{\wedge})\big)d_{b,:,c},&\text{otherwise},\end{cases}$ (37)
[…]
@@
This layer uses the matrix commutator, a natural $\mathrm{Ad}$-equivariant primitive, to create learnable interactions between channels. Let the input be a batch of features represented by their vector coordinates, $x\in\mathbb{R}^{B\times K\times C_{\mathrm{in}}}$.
First, two independent linear maps with weights $W_{a},W_{b}\in\mathbb{R}^{C_{\mathrm{in}}\times C_{\mathrm{out}}}$ produce intermediate tensors $u=xW_{a}$ and $v=xW_{b}$. The Lie bracket is then computed channel-wise between the corresponding feature vectors of $u$ and $v$ to produce an update tensor $\Delta x\in\mathbb{R}^{B\times K\times C_{\mathrm{out}}}$:
$(\Delta x)_{b,:,c^{\prime}}=\left[(u_{b,:,c^{\prime}})^{\wedge},(v_{b,:,c^{\prime}})^{\wedge}\right]^{\vee}.$ (38)
This update is added to the input via a residual connection (requiring $C_{\mathrm{in}}=C_{\mathrm{out}}$):
@@
Each step in this process is equivariant under the adjoint action, making the entire block exactly $\mathrm{Ad}_{g}$-equivariant: $f_{\mathrm{ReLN-Bracket}}(\mathrm{Ad}_{g}x)=\mathrm{Ad}_{g}(f_{\mathrm{ReLN-Bracket}}(x))$ for all $g\in G$.
Appendix F Theoretical Complexity and Scaling Analysis
To address the operational properties of our architecture, we provide an explicit Big-$\mathcal{O}$ complexity analysis connecting the feature channel width $C$ and the Lie algebra matrix dimension $n$. We further validate this analysis with empirical benchmarks.
1. Theoretical Scaling with Matrix Dimension $n$.
We fix the total number of scalars per token to $C$ to ensure a fair comparison. The effective channel width $M$ is scaled such that $M\times K=C$.
  • Standard MLP ($K=1$): Uses full width $M=C$.
  • ReLN ($K=n^{2}$): Operates on $\mathfrak{gl}(n)$ matrices. Effective width $M=C/n^{2}$.
Let $n$ be the dimension of the matrices (e.g., $n=3$ for $\mathfrak{gl}(3)$).
  • Linear Layer: The layer performs $K$ independent matrix multiplications of size $M\times M$.
    $\text{Cost}_{\text{linear}}=K\cdot M^{2}=n^{2}\cdot\left(\frac{C}{n^{2}}\right)^{2}=\frac{C^{2}}{n^{2}}.$ (40)
    The dominant quadratic cost decreases as the matrix dimension $n$ increases. For $n=3$, ReLN reduces the Linear Layer FLOPs by a factor of $1/9$ compared to an MLP ($n=1$).
  • Lie Bracket Nonlinearity: This operation computes [X,Y]=XYYX[X,Y]=XY-YX, involving $n\times n$ matrix multiplications.
    $\text{Cost}_{\text{bracket}}=M\cdot\mathcal{O}(n^{3})=\frac{C}{n^{2}}\cdot\mathcal{O}(n^{3})=\mathcal{O}(C\cdot n).$ (41)
    In deep learning contexts where $C$ is large (e.g., $C\geq 128$), the quadratic term $\mathcal{O}(C^{2}/n^{2})$ dominates the linear term $\mathcal{O}(Cn)$.
2. Layer-wise Complexity Comparison.
Table 8 presents the theoretical breakdown. ReLN-Bracket reduces the total block cost to $\approx 12.5\%$ of the baseline.
Table 8: Theoretical Complexity Analysis. Comparison under a fixed feature budget $C$. By utilizing a higher-dimensional geometric space ($n=3,K=9$), ReLN reduces the effective channel width to $C/9$, resulting in a quadratic reduction in the dominant Linear Layer cost.
Layer Type Model Eff. Channels ($M$) Parameters FLOPs Relative Cost
Linear Layer MLP ($n=1$) $C$ $C^{2}$ $C^{2}$ $1.00\times$
VN ($n\approx 1.7$) $C/3$ $0.33C^{2}$ $C^{2} 0.33×0.33\times$
ReLN ($n=3$) $\mathbf{C/9}$ $\mathbf{0.11C^{2}}$ $\mathbf{0.11C^{2}}$ $\mathbf{0.11\times}$
Nonlinearity ReLU (MLP) $C$ 0 $\mathcal{O}(C)$ $\approx 0.00\times$
VN-ReLU $C/3$ $0.11C^{2}$ $C^{2} 0.11×0.11\times$
ReLN-ReLU $C/9$ $01C^{2} 0.0$$1C^{2} 0.0$$01\times$
ReLN-Bracket $\mathbf{C/9}$ $\mathbf{0.025C^{2}}$ $\mathbf{0.025C^{2}}$ $\mathbf{0.03\times}$
Appendix G Experimental Details
@@
G.1 Model Architectures and Implementation Details
@@
For the Top-Tagging task, our model is a modification of the LorentzNet architecture (Gong et al., 2022). We adapt its Lorentz Group Equivariant Blocks (LGEBs) by replacing the invariant feature computation with our proposed bilinear form. A detailed description of the architecture, our modifications, and training protocol is provided in Appendix H.
@@
In this task, we compare our ReLN model against two baseline families: a non-equivariant 1D ResNet and an equivariant Vector Neurons (VN) model. The specific implementation details and architectural choices for each model are provided next in Appendix I.
G.2 Platonic Solid Classification on $\mathfrak{sl}(3)$
@@
Appendix H Top Tagging Experiment: Framework, Proof, and Implementation
This appendix provides details for our Lorentz-equivariant jet tagging study. We first summarize the task and results, then present the Lorentz-compatible embedding and its equivariance proof, and finally describe the model and training protocol.
H.1 Particle Physics with Lorentz Group $\mathrm{SO}^{+}(1,3)$ Equivariance
We evaluate on the Top-Tagging benchmark (Kasieczka et al., 2019), which classifies particle jets originating from top quarks against a Quantum Chromodynamics background. Since relativistic collisions respect spacetime symmetries, the model should be equivariant under the proper orthochronous Lorentz group $\mathrm{SO}^{+}(1,3)$. To express this as an adjoint action within our Lie-algebraic framework, we embed each four-momentum $p\in\mathbb{R}^{1,3}$ into $\mathfrak{gl}(5)$ via
$\varphi(p)\;=\;\begin{bmatrix}0_{4\times 4}&p\\ p^{\top}\eta&0\end{bmatrix},\qquad\eta=\mathrm{diag}(-1,1,1,1),$ (42)
and show in Appendix H.2 that $\varphi(\Lambda p)=\mathrm{Ad}_{\mathrm{diag}(\Lambda,1)}\big(\varphi(p)\big)$ for $\Lambda\in\mathrm{SO}^{+}(1,3)$.
For comparison, we modify LorentzNet by replacing its invariant feature computation with our bilinear-form construction and additionally report a parameter-matched LorentzNet baseline. Table 11 shows that ReLN achieves performance comparable to the parameter-matched LorentzNet across Acc./AUC and Rej@30%, with differences within the reported uncertainty (i.e., overlapping standard deviations).
Table 11: Comparison of performance on the Top-Tagging dataset. ReLN results include both the parameter-efficient version (84k) and the parameter-matched version (224k). Rej@30% denotes background rejection at 30% signal efficiency (higher is better). Benchmark scores are as reported in the original publications.
Architecture #Params Acc. AUC Rej@30% Reference
PELICAN 45k 0.943 0.987 $2289\pm 204$ Bogatskiy et al. (2022)
LorentzNet (orig.) 224k 0.942 0.987 $2195\pm 173$ Gong et al. (2022)
LorentzNet (reprod.) 84k 0.942 0.987 $1821\pm 94$ Our reprod.
LGN 4.5k 0.929 0.964 $435\pm 95$ Bogatskiy et al. (2020)
BIP 4k 0.931 0.981 $853\pm 68$ Munoz et al. (2022)
partT 2.14M 0.940 0.986 $1602\pm 81$ Qu et al. (2022)
ParticleNet 498k 0.938 0.985 $1298\pm 46$ Qu and Gouskos (2020)
EFN 82k 0.927 0.979 $633\pm 31$ Komiske et al. (2019)
TopoDNN 59k 0.916 0.972 $295\pm 5$ Pearkes et al. (2017)
LorentzMACE 228k 0.942 0.987 $1935\pm 85$ Batatia et al. (2023)
CGENN 321K 0.942 0.987 $2172$ Ruhe et al. (2023)
Lie Neurons 224k 0.941 0.985 $1655\pm 73$ Our reprod.
ReLN (Ours) 224k 0.942 0.987 $2201\pm 101$
ReLN (Ours) 84k 0.942 0.987 $1979\pm 87$
H.2 Group Action Equivariance via Embedding Map
To process four-momenta within our Lie-algebraic framework, we require an embedding that translates the Lorentz action into an adjoint action on a matrix space.
Definition H.1 (Lorentz-Compatible Embedding).
Given $p\in\mathbb{R}^{4}$ and $\eta=\operatorname{diag}(-1,1,1,1)$, define